OpenStudy (anonymous):

Tommy and Sally are arguing over the meanings of x^4 and x^-4. Tommy argues that x^4 is the reciprocal of x^-4. Sally insists that x^-4 is the reciprocal of x^4. Who is correct? A. Tommy is correct. B. Sally is correct. C. Both Tommy and Sally are correct. D. Neither Tommy nor Sally is correct.

OpenStudy (anonymous):

x^4 and x^(-4) are reciprocals of each other. Note: Two numbers are the reciprocals of each other if their product is 1. (x^4) (x^(-4)) = x^0 = 1. @loudestlucid